Which of the following quadratic equation have repeated roots-
1. X square+6x+9=0
2. Z square-6z+6=0
3. 2y square+15y+30=0
4. 9x square-12x+4=0
5. 16y square-40y+25=0
6. 9z square-6z+4=0
Please I want help in this...

Answer by Alan3354(69443)

Which of the following quadratic equation have repeated roots-
1. X square+6x+9=0
2. Z square-6z+6=0
3. 2y square+15y+30=0
4. 9x square-12x+4=0
5. 16y square-40y+25=0
6. 9z square-6z+4=0
-----------
They're all in the form ax^2 + bx + c = 0
---
Find the discriminant.
Disc = b^2 - 4ac
1. X^2 + 6x + 9=0
a = 1, b = 6, c = 9
Disc = 6^2 - 4*1*9 = 36 - 36 = 0
--------------
If the Disc = 0, the 2 roots are equal.
===========================
Do the others the same way.